Solve the following thermochemical reactions:

2 HCl + Pb


PbCl2 + H2


H = -1861 J/mole

What is the enthalpy if there was 3.86g of PbCl2?


1
Expert's answer
2022-06-15T14:20:04-0400

2HCl + Pb = PbCl2 + H2 ∆H = -1861 J/mol


M (PbCl2) = 278 g/mol


278 g PbCl2 — ∆H = -1861 J

3.86 g PbCl2 — ∆H = X


X = 3.86 * (-1861) / 278 = -25.84



Answer: ∆H = -25.84 J